OpenStudy (anonymous):

Triangle ABC is similar to triangle FGH. Angle A measures 42 degrees. Angle B measures 19 degrees. Find the measure of angle H

OpenStudy (anonymous):

Angle total = 180 = 42 + 19 + angle H => angle H = 180 - 42 - 19 = 119